# On the Noisy Road to Open Quantum Dynamics: The Place of Stochastic Hamiltonians

### v1

@@ Abstract
Stochastic evolution underpins several approaches to the dynamics of open quantum systems, such as random modulation of Hamiltonian parameters, the stochastic Schrödinger equation (SSE), and the stochastic Liouville equation (SLE). In a stochastic formulation, the open-system problem is reduced from a coupled system-environment dynamics to an effective system-only description, with dissipative evolution recovered by ensemble averaging. In this work, we aim at a self-contained and accessible presentation of these approaches to further elaborate on their common roots in essential concepts of stochastic calculus and to delineate the conditions under which they are equivalent. We also discuss how different formulations naturally lead to different numerical time-integration schemes, better suited for either classical simulation platforms, based on finite-difference approximations, or quantum algorithms, that employ random unitary maps. Our analysis supplies a unified perspective and actionable recipes for classical and quantum implementations of stochastic evolution in the simulation of open quantum systems.
@@ I Introduction
Quantum systems are never fully isolated. Molecules and nano-structures, both natural and artificial, are embedded in some medium, such as solid-state matrices or fluids. Even almost-isolated systems in high vacuum, such as the neutral atoms and trapped ions used in quantum computing architectures, interact with the control apparatus and external fields. The surroundings and all the secondary degrees of freedom are part of the so-called environment of the system. The unavoidable interactions with the environment influence the expected ideal behavior of quantum systems, inducing decoherence and dissipation [petruccione2002, Vacchini2024OpenTheory], as they cannot be considered closed systems any longer. An effective treatment of open quantum systems then requires shifting from the unitary evolution of a wavefunction to a density matrix description, whose evolution does not follow a universal equation of motion but instead depends on the approximations and modeling strategies that one employs.
@@
In this work, we start from the dynamics induced by a stochastic Hamiltonian [Kubo1963StochasticEquations, Haken1973AnMotion, Hasegawa1980StochasticProcesses], as it provides a particularly direct and versatile route. In this framework, the effective Hamiltonian $H^{\mathrm{eff}}_{t}$ is composed of the bare system Hamiltonian $H$ plus a randomly fluctuating term $H_{t}^{\mathrm{fluct}}$, so that we can write
@@
Direct introduction of random fluctuations in the Hamiltonians is not the only way of introducing stochastic dynamics in the quantum evolution. The stochastic Schrödinger equation (SSE) is a more general formalism in which the pure state of an open quantum system evolves according to a stochastic differential equation driven by noise terms [Gisin1984QuantumProcesses, Barchielli1991MeasurementsMechanics, Gisin1992TheSystems, Carmichael1993An18]. These noise terms were interpreted in connection with specific measurement protocols, “quantum jumps” or continuous measurement back-action [petruccione2002, Jacobs2006AMeasurement, Barchielli2009Quantum782]. Ultimately, the SSE is constructed so that the ensemble average over trajectories reproduces the deterministic evolution dictated by a master equation (in Lindblad form or non-Markovian extensions thereof [Barchielli2009Quantum782, Diosi1997TheSystems, Breuer1999StochasticEquations, Semina2014StochasticCases, Li2021MarkovianEquations, Gasbarri2018StochasticMaps]).
Renewed interest in stochastic methods in quantum dynamics has recently emerged in diverse contexts, ranging from molecular modeling [Coccia2018ProbingSystems, DallOsto2024StochasticSystems] to quantum technologies [Huang2020QubitNoise, DeKeijzer2025QubitNoise]. In this paper, we aim for a solid and clear analysis of these methods by uncovering and elaborating on their common roots in the basic concepts of stochastic calculus. We aim at offering a unified perspective that bridges methodological developments using stochastic Hamiltonians across chemical physics and quantum information science.
@@
The stochastic formulation of the open-system dynamics replaces an explicit system–environment evolution with an effective system-only description, drastically reducing complexity. It also offers an efficient platform for the simulation of many-body systems (spin glasses, molecules, nanostructures), where computing the density-matrix evolution is often prohibitive due to the quadratic scaling of the state dimensions, compared to the linear scaling of the wavefunction. However, the shift from a deterministic to a stochastic dynamics affects both the formal apparatus and the interpretive framework.
@@
In what follows, we present a self-contained analysis of stochastic processes in open quantum dynamics. Our primary goal is to clarify the interrelations among stochastic Hamiltonians, the stochastic Schrödinger equation (SSE), and the stochastic Liouville equation (SLE), and to delineate the conditions under which they are equivalent. To this end, we will make use of the two main interpretations of stochastic calculus, named after Itō and Stratonovich, presented in Section II.2, and point out explicitly the difference between a stochastic process and a noise. We further show how each formulation naturally maps to distinct numerical workflows, amenable to efficient classical trajectory integration or to quantum-algorithmic implementations. Figure 1 illustrates graphically the results presented in Sections III, LABEL:sec:ito and LABEL:sec:rode. First, in Section III, we discuss the lower blue pathway. Equation 6 is interpreted in the Stratonovich formalism as a viable way to implement quantum simulations, leading directly to the composition of universal gates in a quantum computing architecture. In LABEL:sec:ito, we follow the upper green branch of the schematic map and adopt the Itō interpretation, showing that it is the correct framework to obtain valid quantum master equations (QME). We also discuss how to write stochastic Schrödinger equations (SSE), and clarify how a stochastic Hamiltonian model is a special case of the SSE subject to more stringent constraints. In addition, we show that the Itō form of the evolution equations is easily treated with classical efficient time-integration schemes. A particular class of equations that can be cast indifferently in both Itō and Stratonovich frameworks is briefly discussed in LABEL:sec:rode. Finally, in LABEL:sec:sle, we show how we can work directly with stochastic density matrices instead of wavefunctions, moving back to the framework of earlier seminal works that employed stochastic Liouville equations [Kubo1963StochasticEquations, Haken1973AnMotion]. This allows for the inclusion of various effects, such as different dissipators and measurement apparatuses.
@@
Figure 1: A schematic map of the two mathematical frameworks and relative workflows for the stochastic Hamiltonian applied to a wavefunction dynamics, based on the Itō (upper green branch) and Stratonovich (lower blue branch) interpretations. The Stratonovich pathway connects stochastic Hamiltonians to direct simulations of the density matrix dynamics through random unitary gates, while the Itō pathway provides the correct formulation for analytically obtaining quantum master equations. The random ordinary differential equation (RODE) case appears as a special limit when dealing with continuous stochastic processes as Hamiltonian fluctuations.
@@ II.1 Modeling of the stochastic Hamiltonian
where $\hat{S}$ acts on the system and $\hat{B}$ is a Hermitian operator acting on the environment. To confine the description within the Hilbert space of the system only, the environment operator $\hat{B}$ must be replaced with a function $B_{t}$. This is the first step of numerous embedding schemes, where the environment is reflected in a possibly time-dependent potential. Notice that when such a potential has a stochastic character, the expected change of the purity of the system state, typical of open quantum systems, can be recovered by averaging over different trajectories [Kubo1966TheTheorem, Gaspard1999Non-MarkovianEquation]. In some cases, the stochastic field can be considered as a surrogate environment, originating from the coupling Hamiltonian and the state of the original microscopic environment [Szankowski2020NoiseDynamics, Szankowski2021MeasuringNoise, Gu2019WhenNoise]. A prominent example used in triplet exciton dynamics is the replacement of a wide-bandwidth bosonic bath in the high-temperature limit with Gaussian white noise, which acts as a genuine surrogate field [Haken1973AnMotion].
The interaction Hamiltonian, now acting on the system subspace only, becomes
@@
The reduced density matrix of the open system is obtained as the mean density matrix over all the possible trajectories, that is
@@
where $\rho^{\mathrm{trj}}_{t}$ is the density matrix of the system along a trajectory. In practice, the average on the ensemble $\Omega$ of trajectories $\omega\in\Omega$ is approximated by a finite sum as
@@ II.2 Stochastic calculus tools
Introducing stochastic potentials to model random contributions to a system’s dynamics naturally calls for the use of stochastic calculus tools. Within this section, we provide a brief review of essential concepts. We recall the differences between the approaches by Itō and Stratonovich to stochastic calculus as they will be central to appreciate the application-oriented analysis of the following sections.
@@
where $\mathop{}\!\mathrm{d}{X_{t}}$ is the stochastic differential of the process $X_{t}$, $f(t,X_{t})$ and $g(t,X_{t})$ are either functionals of the process or simple functions of time, $W_{t}$ is a Brownian motion (Wiener process), and $\mathop{}\!\mathrm{d}{W_{t}}$ its differential with scaling $\mathop{}\!\mathrm{d}{W_{t}}\sim\sqrt{\mathop{}\!\mathrm{d}{t}}$. Indeed, this type of notation is needed because the time derivative of the Wiener process is not a well-defined function; hence, the same is true for the time derivative of the generic process $X_{t}$.
@@
Itō’s integration is identified by $\lambda\!=\!0$, which corresponds to setting $\tau_{k}\!=t_{k}$, the left-hand point of each subinterval. This choice means that we do not need previous knowledge of the function at future values, in agreement with the idea that the stochastic potential provides unpredictable random perturbations to the system. In this interpretation of the stochastic differentials, the product rule for the differentials is not the ordinary one, but Itō’s product rule. For two differentials of the form $\mathop{}\!\mathrm{d}X_{i}=F_{i}\mathop{}\!\mathrm{d}t+G_{i}\mathop{}\!\mathrm{d}W_{t}$, the differential of the product is
@@
where the usefulness of restoring the basic rules of differential calculus justifies the tricky anticipatory nature of this interpretation. A fundamental point now is the definition of the $\circ$ symbol as the symmetric Q-multiplication [Ito1975StochasticCalculus] which, for two time-dependent random variables $X_{t}$ and $Y_{t}$, is
@@
with the differential on the r.h.s. now expressed in the Itō interpretation. Then, by combinations of eqs. 16 and 15 we can transform a Stratonovich SDE into an Itō SDE.
Hereforth, we define a stochastic process as the integrated form of any Itō stochastic differential equation, $X=\int\mathop{}\!\mathrm{d}{X_{t}}$, that is the solution of eq. 10. The term noise corresponds instead to the formal time derivative of the stochastic process in a distributional sense, $\dot{X_{t}}=\mathop{}\!\mathrm{d}{X_{t}}/\mathop{}\!\mathrm{d}{t}$. Therefore, in the SDE formulation we define noise-inducing the stochastic differential $\mathop{}\!\mathrm{d}{X_{t}}$ itself. This definition can be understood by formally dividing eq. 9 by the time differential $\mathop{}\!\mathrm{d}{t}$ and obtaining a Langevin-like form, where the presence of a noisy drive is evident in the equation of motion for the process.

### v2

@@ Abstract
Stochastic evolution underpins several approaches to the dynamics of open quantum systems, such as random modulation of Hamiltonian parameters, the stochastic Schrödinger equation (SSE), and the stochastic Liouville equation (SLE). These approaches replace the explicit system–environment coupling with an effective system-only dynamics, where dissipative behavior emerges from ensemble averaging. Stochastic Hamiltonians, in particular, have long served as phenomenological tools in physical chemistry to include environmental effects without recourse to an explicit microscopic derivation. In this work, we aim at a self-contained and accessible presentation of these approaches to further elaborate on their common roots in essential concepts of stochastic calculus and to delineate the conditions under which they are equivalent. We also discuss how different formulations naturally lead to different numerical time-integration schemes, better suited for either classical simulation platforms, based on finite-difference approximations, or quantum algorithms, that employ random unitary maps. Our analysis aims at providing a unified perspective and actionable recipes for classical and quantum implementations of stochastic evolution in the simulation of open quantum systems.
@@ I Introduction
Quantum systems are never fully isolated. Molecules and nano-structures, both natural and artificial, are embedded in some medium, such as solid-state matrices or fluids. Even almost-isolated systems in high vacuum, such as the neutral atoms and trapped ions used in quantum computing architectures, interact with the control apparatus and external fields. The surroundings and all the secondary degrees of freedom are part of the so-called environment of the system. The unavoidable interactions with the environment influence the expected ideal behavior of quantum systems, inducing decoherence and dissipation [petruccione2002, Vacchini2024OpenTheory]. An effective treatment of open quantum systems then requires shifting from the unitary evolution of a wavefunction to a density matrix description, whose evolution does not follow a universal equation of motion but instead depends on the approximations and modeling strategies that one employs.
@@
In this work, we start from the dynamics induced by a stochastic Hamiltonian [Kubo1963StochasticEquations, Haken1973AnMotion, Hasegawa1980StochasticProcesses], as it provides a particularly direct and versatile route to model environmental effects without the need for a complete microscopic model of the underlying system-environment dynamics. In this framework, the effective Hamiltonian $H^{\mathrm{eff}}_{t}$ is composed of the bare system Hamiltonian $H$ plus a randomly fluctuating term $H_{t}^{\mathrm{fluct}}$, so that we can write
@@
Direct introduction of random fluctuations in the Hamiltonian is not the only way of introducing stochastic dynamics in the quantum evolution. The stochastic Schrödinger equation (SSE) is a more general formalism in which the pure state of an open quantum system evolves according to a stochastic differential equation driven by noise terms [Gisin1984QuantumProcesses, Barchielli1991MeasurementsMechanics, Diosi1988ContinuousFormalism, Diosi1988LocalizedEquation, Gisin1992TheSystems, Carmichael1993An18]. These noise terms were interpreted in connection with specific measurement protocols, “quantum jumps” or continuous measurement back-action [petruccione2002, Jacobs2006AMeasurement, Barchielli2009Quantum782]. Ultimately, the SSE is constructed so that the ensemble average over trajectories reproduces the deterministic evolution dictated by a master equation (in Lindblad form or non-Markovian extensions thereof [Barchielli2009Quantum782, Diosi1997TheSystems, Breuer1999StochasticEquations, Semina2014StochasticCases, Li2021MarkovianEquations, Gasbarri2018StochasticMaps]).
@@
Renewed interest in stochastic methods in quantum dynamics has recently emerged in diverse contexts, ranging from molecular modeling [Coccia2018ProbingSystems, DallOsto2024StochasticSystems] to quantum technologies [Huang2020QubitNoise, DeKeijzer2025QubitNoise]. The stochastic formulation of the open-system dynamics replaces an explicit system–environment evolution with an effective system-only description, drastically reducing complexity. It also offers an efficient platform for the simulation of many-body systems (spin glasses, molecules, nanostructures), where computing the density-matrix evolution is often prohibitive due to the quadratic scaling of the state dimensions, compared to the linear scaling of the wavefunction. However, the shift from a deterministic to a stochastic dynamics affects both the formal apparatus and the interpretive framework.
In this paper, we aim to provide a clear analysis of these methods by uncovering and elaborating on their common roots in the basic concepts of stochastic calculus. We wish to present a unified perspective that bridges methodological developments using stochastic Hamiltonians across chemical physics and quantum information science. This work is not intended to provide a comprehensive review of all the stochastic approaches to open quantum systems; rather, it assembles selected tools and results known from specific applications and draws explicit connections among them within a single stochastic-calculus framework. Our main interest is in dynamics governed by stochastic Hamiltonians (SH), which we identify as a constrained instance of the stochastic Schrödinger equation (SSE) for pure states and of the stochastic Liouville equation (SLE) for generally mixed states. We trace how different choices in the formalism determine implementation pathways: numerical quadrature of stochastic trajectories versus random-unitary maps, the latter being well-suited for integration into quantum algorithms. In this sense, the contribution of this work is a self-contained synthesis that clarifies foundations, equivalences, and computational consequences, yielding a practical toolkit for simulating open-system dynamics with stochastic Hamiltonians.
@@
To this end, we will make use of the two main interpretations of stochastic calculus, named after Itō and Stratonovich, presented in Section II.2, and point out explicitly the difference between a stochastic process and a noise. We further show how each formulation naturally maps to distinct numerical workflows, amenable to efficient classical trajectory integration or to quantum-algorithmic implementations. Figure 1 illustrates graphically the results presented in Sections III, LABEL:sec:ito and LABEL:sec:rode. First, in Section III, we discuss the lower blue pathway. Equation 6 is interpreted in the Stratonovich formalism as a viable way to implement quantum simulations, leading directly to the composition of universal gates in a quantum computing architecture. In LABEL:sec:ito, we follow the upper green branch of the schematic map and adopt the Itō interpretation, showing that it is the correct framework to obtain valid quantum master equations (QME). We also discuss how to write stochastic Schrödinger equations (SSE), and clarify how a stochastic Hamiltonian model is a special case of the SSE subject to more stringent constraints. In addition, we show that the Itō form of the evolution equations is easily treated with classical efficient time-integration schemes. A particular class of equations that can be cast indifferently in both Itō and Stratonovich frameworks is briefly discussed in LABEL:sec:rode. Finally, in LABEL:sec:sle, we show how we can work directly with stochastic density matrices instead of wavefunctions, moving back to the framework of earlier seminal works that employed stochastic Liouville equations [Kubo1963StochasticEquations, Haken1973AnMotion]. This allows for the inclusion of various effects, such as different dissipators and measurement apparatuses.
@@
Figure 1: A schematic map of the two mathematical frameworks and relative workflows for the stochastic Hamiltonian applied to a wavefunction dynamics, based on the Itō (upper green branch) and Stratonovich (lower blue branch) interpretations. The Stratonovich pathway connects stochastic Hamiltonians to direct simulations of the density matrix dynamics through random unitary gates, while the Itō pathway provides the correct formulation for analytically obtaining quantum master equations (QME). The random ordinary differential equation (RODE) case appears as a special limit when dealing with continuous stochastic processes as Hamiltonian fluctuations.
@@ II.1 Modeling of the stochastic Hamiltonian
where $\hat{S}$ acts on the system and $\hat{B}$ is a Hermitian operator acting on the environment. To confine the description within the Hilbert space of the system only, the environment operator $\hat{B}$ must be replaced with a function $B_{t}$. This is the first step of numerous embedding schemes, where the environment is reflected in a possibly time-dependent potential. Notice that when such a potential has a stochastic character, the expected change of the purity of the system state, typical of open quantum systems, can be recovered by averaging over different trajectories [Kubo1966TheTheorem, Gaspard1999Non-MarkovianEquation].
In some cases, the stochastic field can be considered as a surrogate environment, originating from the coupling Hamiltonian and the state of the original microscopic environment [Szankowski2020NoiseDynamics, Szankowski2021MeasuringNoise, Gu2019WhenNoise]. A prominent example used in triplet exciton dynamics is the replacement of a wide-bandwidth bosonic bath in the high-temperature limit with Gaussian white noise, which acts as a genuine surrogate field [Haken1973AnMotion]. More commonly, the use of stochastic fluctuations in the Hamiltonian is postulated to represent the dominant environmental influence, with no reference to a microscopically defined environment. This phenomenological approach is conceptually different from the exact treatment of open quantum dynamics, where the elimination of the environment’s degrees of freedom leads to specific stochastic forms like those originating from the Feynman-Vernon (path-integral) formalism [Feynman1963TheSystem, Grabert1988QuantumApproach], non-Markovian quantum state diffusion (NMQSD) [Diosi1997TheSystems, Diosi1998Non-MarkovianDiffusion, Stockburger2001Non-MarkovianDiffusion], the hierarchy of pure states (HOPS) [Suess2014HierarchyDynamics], the HEOM approach[Tanimura1989TimeBath, Tanimura1990NonperturbativeBath, Tanimura2020NumericallyHEOM], the generalized Langevin equations (GLE) [Stella2014GeneralizedSystems, Ness2015ApplicationsBaths, Matos2020EfficientSystems] and others [Stockburger2002ExactDissipation, McCaul2017Partition-freeEquation, Magazzu2022Feynman-VernonStudy].
The fluctuating interaction Hamiltonian, now acting on the system subspace only, becomes
@@
The reduced density matrix of the open system is obtained as the averaged density matrix over all the possible trajectories, that is
@@
where $\rho^{\mathrm{trj}}_{t}$ is the density matrix of the system along a trajectory and $\mathbb{E}[\cdot]$ indicates the expected value of the argument, obtained as the ensemble average. In any practical case, the average over the ensemble of all possible trajectories $\omega$ is approximated by a finite averaging
@@
We remark that the averaged density matrix, eq. 7, does not describe a pure state, but rather the mixed state of an ensemble where dissipative effects stem from the fluctuations induced by the stochastic component. Importantly, the average maps are completely positive and trace-preserving (CPT) [Gasbarri2018StochasticMaps] by construction, as the mean dynamics derives from the average of pure quantum states.
[…]
Figure 2: Example of the effect of a stochastic fluctuation at a time $t$ on the Hamiltonian energy levels in site basis for the first exciton manifold of a system of two degenerate sites, i.e. $\varepsilon_{0}=\varepsilon_{1}$, indicated by the central dotted line. The vertical blue lines indicate the variance of the stochastic energy fluctuation, the red lines the instantaneous energy of each level at a time $t$, namely $\varepsilon_{i}+\delta\varepsilon_{i}(t)$. The two levels are coupled by a term of constant $\Omega$, shown as a black line.
To exemplify the formalisms introduced above and the further development in the following sections, we will consider examples of quantum two-level systems. In what follows, we recall the stochastic Hamiltonian formulation of coherence decay between two excited states, originally proposed in the context of triplet exciton dynamics by Haken and Strobl [Haken1973AnMotion].
Example 1 (Haken-Strobl) Let the system basis be composed of two sites $\{|0\rangle,|1\rangle\}$, with energy $\varepsilon_{0}$ and $\varepsilon_{1}$, respectively, and let them interact by a coupling $\Omega$ that coherently switches between the sites. The total Hamiltonian, $H$ in eq. (1), is then
$H=\varepsilon_{0}|0\rangle\langle 0|+\varepsilon_{1}|1\rangle\langle 1|+\Omega\big(|0\rangle\langle 1|+|1\rangle\langle 0|\big)\,.$ (9)
A stochastic fluctuation, modeled as white noise (see LABEL:app:whitenoise), is applied to the site energies to replicate the effect of a rapid Markovian environment, fig. 2. We can therefore write the fluctuation Hamiltonian as
$\hat{H}_{t}^{\mathrm{fluct}}=\delta\varepsilon_{0}(t)|0\rangle\langle 0|+\delta\varepsilon_{1}(t)|1\rangle\langle 1|\,.$ (10)
From a comparison with eq. (1), we can identify two noise variables $Z_{0}(t)=\delta\varepsilon_{0}(t)$ and $Z_{1}(t)=\delta\varepsilon_{1}(t)$, and the two associated operators $R_{0}=|0\rangle\langle 0|$ and $R_{1}=|1\rangle\langle 1|$. This setting reflects a Haken-Strobl model [Haken1973AnMotion], the prototypical description of site dephasing in a quantum system. The trajectory dynamics are computed according to eq. 6, the formal explanation for this choice will become clear in Section III, and the mean dynamics is obtained as the average over trajectories, see eq. 8. It is well known (see LABEL:sec:ito for the details) that the average dynamics of this model is exactly the solution of a Lindblad equation, reported in LABEL:app:HS_Lindblad. In fig. 3, a swarm of trajectories is depicted, showing the random character of each realization, and the effect of the different sizes of the sample space is highlighted in the two panels by comparison with the exact solution of the associated Lindblad master equation. Note that, differently from the average behavior, each trajectory maintains the coherent (oscillating) character typical of pure states. The system parameters used for the example are reported in LABEL:app:HS_Lindblad.
Figure 3: Stochastic trajectories and average dynamics of the excited-state population $\rho^{(11)}_{t}=\langle 1|\rho_{t}|1\rangle$ for the Haken–Strobl model presented in the text, computed with the stochastic Hamiltonian approach. (a) Individual realizations of a small ensemble of $10^{2}$ trajectories (grey lines in transparency), one single trajectory highlighted (blue line), and their average (green line). In panel (b), the average over $10^{3}$ trajectories (black line) is compared with the exact solution of the corresponding Lindblad master equation (red dashed line), showing agreement between the stochastic and the master equation descriptions.
@@ II.2 Stochastic calculus tools
Introducing stochastic potentials to model random contributions to a system’s dynamics naturally calls for the use of stochastic calculus tools. Within this section, we provide a brief review of essential concepts. We recall the differences between the approaches by Itō [Ito1944StochasticIntegral] and Stratonovich [Stratonovich1966AEquations] to stochastic calculus as they will be central to appreciate the application-oriented analysis of the following sections.
@@
where $\mathop{}\!\mathrm{d}{X_{t}}$ is the stochastic differential of the process $X_{t}$, $f(t,X_{t})$ and $g(t,X_{t})$ are either functionals of the process or simple functions of time, $W_{t}$ is a Brownian motion (Wiener process), and $\mathop{}\!\mathrm{d}{W_{t}}$ its differential with scaling $\mathop{}\!\mathrm{d}{W_{t}}\sim\sqrt{\mathop{}\!\mathrm{d}{t}}$, see LABEL:app:whitenoise for more details. Indeed, this type of notation is needed because the time derivative of the Wiener process is not a well-defined function; hence, the same is true for the time derivative of the generic process $X_{t}$.
@@
Itō’s integration is identified by $\lambda\!=\!0$, which corresponds to setting $\tau_{k}\!=t_{k}$, the left-hand point of each subinterval. This choice means that we do not need previous knowledge of the function at future values, in agreement with the idea that the stochastic potential provides unpredictable random perturbations to the system. In this interpretation of the stochastic differentials, the product rule for the differentials is not the ordinary one, but Itō’s product rule: for two differentials of the form $\mathop{}\!\mathrm{d}X_{i}=F_{i}\mathop{}\!\mathrm{d}t+G_{i}\mathop{}\!\mathrm{d}W_{t}$, the differential of the product is
@@
where the usefulness of restoring the basic rules of differential calculus justifies the tricky anticipatory nature of this interpretation. A fundamental point now is the definition of the $\circ$ symbol we used to indicate the Stratonovich differential as an actual multiplication, known as the symmetric Q-multiplication [Ito1975StochasticCalculus] which, for two time-dependent random variables $X_{t}$ and $Y_{t}$, is
@@
with the differential on the r.h.s. now expressed in the Itō interpretation. Then, by combinations, and possibly iteration, of eq. 17 we can transform a Stratonovich SDE into an Itō SDE.
Hereforth, we define a stochastic process as the integrated form of any Itō stochastic differential equation, $X=\int\mathop{}\!\mathrm{d}{X_{t}}$, that is the solution of eq. 12. The term noise corresponds instead to the formal time derivative of the stochastic process in a distributional sense, $\dot{X_{t}}=\mathop{}\!\mathrm{d}{X_{t}}/\mathop{}\!\mathrm{d}{t}$. Therefore, in the SDE formulation, we define noise-inducing the stochastic differential $\mathop{}\!\mathrm{d}{X_{t}}$ itself. This definition can be understood by formally dividing eq. 11 by the time differential $\mathop{}\!\mathrm{d}{t}$ and obtaining a Langevin-like form, where the presence of a noisy drive is evident in the equation of motion for the process.